A wagon is moving with a velocity of -2.0 m/s while having a 20 kg mass on top of it. There is a force applied to the wagon over a distance of 1.25 m. It then comes to a stop with a velocity of 0m/s. Please answer the questions. Show all work.

1. What is the total work done on the wagon?

2. What was the average force applied to the box to bring it to a halt?
Answer #1

a) for the work done on the wagon

Work done on the wagon = change in kinetic energy of wagon

Work done on the wagon = 0.50 * 20 * (0^2 - 2^2)

Work done on the wagon = -40 J

the total work done on the wagon is -40 J

b) NOw, for the average force

Work done on the wagon = F * d * cos(theta)

-40 = F * (1.25) * cos(180)

F = 32 N

the average force applied is 32 N